OpenStudy (brittanydosey):

You invest $5,175.00. It increases 19% the first year, and loses 15% of its value the second year. What is your gain compared to your original investment?

OpenStudy (brittanydosey):

@killerdime20

OpenStudy (brittanydosey):

@hysusonic

OpenStudy (brittanydosey):

i had a bit of trouble with this one

OpenStudy (anonymous):

Each stock would be worth 5419.26 after those 2 years. 5419.26-5175.00=244.26

OpenStudy (anonymous):

it increases 19% the first year 5175 + 0.19(5175) = 5175 + 983.25 = 6158.25 2nd yr...loses 15% 6158.25 - 0.15(6158.25) = 6158.25 - 923.74 = 5234.51 gain - original investment = 5234.51 - 5175 = 59.51

OpenStudy (jdoe0001):

get the 19% gain first \(\bf {\color{brown}{ 19}}\%\ of\ 5,175.00 \implies \cfrac{{\color{brown}{ 19}}}{100}\cdot 5,175.00\implies {\color{brown}{ 19}}\div 100\cdot 5,175.00\) then get the 17% from that amount add the "gain" to the original amount and subtract the "loss" from that gained new amount that'd be the total amount left after 2 years check the difference

OpenStudy (brittanydosey):

thank you, can you help me on 4 more?

OpenStudy (anonymous):

I can try